# Exact Quench Dynamics from Thermal Pure Quantum States

### v1

@@ Abstract
We present an exact solution for a quantum quench in an integrable system that reveals a new, coherent pathway to thermalization. In chaotic systems, thermalization is understood via the eigenstate thermalization hypothesis (ETH), which implies that a quench from a thermal pure quantum (TPQ) state should exhibit trivial dynamics. In contrast, integrable systems are expected to retain memory of their initial conditions in a non-thermal generalized Gibbs ensemble (GGE). We show that a quench from a TPQ state in the integrable spin-1/2 XX chain defies both expectations. The system thermalizes to a canonical Gibbs ensemble, yet the entanglement dynamics are highly non-trivial, exhibiting a ”double plateau” structure. We solve this process exactly using techniques in two-dimensional (2D) conformal field theories (CFT), an exact numerical method based on the matrix Riccati equation, and an asymptotically exact quasiparticle picture. We prove that the non-trivial dynamics are the macroscopic signature of the coherent dephasing of the initial state’s anomalous pairing correlations.
@@
A central pillar of modern statistical mechanics is the eigenstate thermalization hypothesis (ETH), which explains how isolated, chaotic quantum systems thermalize [1, 2, 3]. The ETH posits that individual energy eigenstates of chaotic systems are already locally thermal. This is a manifestation of canonical typicality: a generic pure state from a microcanonical energy shell is locally indistinguishable from a thermal ensemble [4]. A direct consequence is that the quench dynamics from a state that is already thermal-like, such as a thermal pure quantum (TPQ) state [5], should be trivial: local observables, including entanglement entropy, should remain constant. In sharp contrast, integrable systems, possessing an extensive number of conservation laws, are known to defy ETH. They retain memory of their initial state and relax to a non-thermal generalized Gibbs ensemble (GGE) [6, 7].
@@
This raises a fundamental question: what the quench dynamics is if an integrable system is prepared in a TPQ state? Does its evolution reflect the underlying memory of integrability, or the thermal nature of its initial state? In this Letter, we provide an exact solution to this problem, uncovering a third dynamical paradigm that is neither the trivial evolution of a chaotic system nor the non-thermal relaxation to a GGE. We study the evolution from a special kind TPQ state $|\Psi_{\beta}\rangle$ in the integrable spin-1/2 XX chain, and the post quench dynamics is also governed by the same Hamiltonian $H$. This TPQ state is constructed by evolving a non-local, maximally entangled crosscap state $|\mathcal{C}\rangle$ [8, 9, 10, 11, 12, 13, 14, 15] in imaginary time: $|\Psi_{\beta}\rangle\propto e^{-\beta H/4}|\mathcal{C}\rangle$[17, 18, 16].
Our central result is that this system thermalize, yet its approach to equilibrium is marked by highly non-trivial, coherent dynamics. We study this problem from three complementary perspectives. First, using techniques in 2D CFT, we derive an analytical formula for the time-dependent entanglement entropy in terms of elliptic theta functions. This is achieved by computing vertex operator correlators on the Klein bottle geometry associated with the state. Second, we provide a rigorous numerical benchmark by deriving the exact initial covariance matrix of the TPQ state by analytically solving the governing matrix Riccati equation, and then evolving it in real time. Our CFT computations and numerical simulations reveal a universal ”double plateau” in the entanglement entropy. This is in stark contrast to the constant entropy expected for a chaotic system under the same protocol. Thirdly, we demonstrate that this entire dynamical profile can be quantitatively explained by the quasiparticle picture [19, 20, 21, 22, 23]. The dynamics are governed by the ballistic transport of pre-existing, antipodally entangled pairs whose properties are determined by $\beta$. The initial decrease in entropy occurs when both members of these non-local pairs enter a subsystem, while the subsequent rise is due to their separation as they traverse in the system. This semiclassical model, which is asymptotically exact in the thermodynamic limit, precisely reproduces the full dynamics.
While it has long been known that integrable CFTs can exhibit thermalization after quenches from boundary states, integrable lattice models almost always relax to GGE [24]. Our work presents a rare example of an integrable lattice model where a pure state evolves coherently to a true Gibbs thermal state, highlighting that thermalization is possible even in discrete integrable systems.
@@
Thermal Pure Quantum State — In 2D CFT, a crosscap state $|\mathcal{C}\rangle$ is a boundary state [25] embodying a non-orientable spacetime, defined by the constraint $(L_{n}-(-1)^{n}\bar{L}_{-n})|\mathcal{C}\rangle=0$ on the Virasoro generators [8]. While abstract, its physical essence is captured on a lattice of length $L$ as an entangled antipodal pair (EAP) state [16]. For a spin-1/2 chain, this is:
@@
The Numerical Benchmark — To validate our analytical CFT predictions, we perform exact numerical simulations on the lattice realization of the theory: the spin-1/2 XX chain at half filling. The Hamiltonian for a system of $L=2N$ sites with periodic boundary conditions is
@@
After a Jordan-Wigner transformation: $\sigma_{j}^{+}=\left(\prod_{k=1}^{j-1}(-\sigma_{k}^{z})\right)c_{j}^{\dagger}$, $\sigma_{j}^{-}=\left(\prod_{k=1}^{j-1}(-\sigma_{k}^{z})\right)c_{j},\sigma_{j}^{z}=2c_{j}^{\dagger}c_{j}-1$. The model is described by a Hamiltonian of non-interacting fermions, $H=-2J\sum_{j=1}^{2N}(c_{j}^{\dagger}c_{j+1}+\text{h.c.})$, with an anti-periodic boundary condition. The single-particle energy dispersion is given by $E(k)=-4J\cos(k)$. The ground state is formed by filling all negative energy states, which corresponds to half-filling ($N_{f}=N$). The low-energy excitations occur near the two Fermi points, $k_{F}=\pm\pi/2$, where the dispersion becomes linear: $E(k_{F}\pm q)\approx\pm(4J)q$. This linear dispersion is the defining characteristic of the massless (1+1)D Dirac fermion, with an emergent ”speed of light” (Fermi velocity) of $v_{F}=4J$ (in units where the lattice spacing $a=1$).
@@
Figure 1: Comparison between the exact numerical results and the quasiparticle prediction for the entanglement entropy $S_{A}(t)$. Solid lines: numerical data; dashed line: quasi-particle prediction. Left: $L=600,l=200,\beta=1$. Right: $L=1000,l=400,\beta=4$.
@@
Mechanism of Thermalization — We now provide the microscopic explanation for the thermalization observed in our integrable system. The process is not driven by chaos, but by the dephasing of the initial state’s quantum coherences, a mechanism that is quantitatively accessible in our model. We show that the stationary state is a GGE that, for this specific initial state, rigorously reduces to the canonical Gibbs ensemble.
A central feature of the initial TPQ state $|\Psi_{\beta}\rangle$ is its non-zero anomalous pairing amplitude, $\phi_{k}(\beta)=\braket{c_{k}c_{-k}}=\frac{-i(-1)^{m}}{2\cosh(E(k)\beta/2)}$. This is a signature of quantum coherence between states with different fermion numbers. The final thermal Gibbs ensemble, however, is an incoherent statistical mixture and must have zero pairing. The mechanism of thermalization must therefore explain how this initial quantum coherence is dynamically lost. We show that this occurs via dephasing, and that the stationary state is a GGE that correctly describes the system after this coherence has vanished. Using the BCS representation of the TPQ state in the thermodynamic limit (11), the time-evolved state is simply
$|\Psi_{\beta}(t)\rangle=\prod_{k>0}\left(u_{k}(\beta)+v_{k}(\beta)e^{-2iE(k)t}c_{k}^{\dagger}c_{-k}^{\dagger}\right)|0\rangle\,.$ (15)
From the equation above, it’s easy to see that the mode occupations $n_{k}$ is constants in time, while he pairing amplitude has a simple oscillatory time dependence $\phi_{k}(t,\beta)=\phi_{k}(0,\beta)e^{-i2E(k)t}$. The long-time average of this oscillating quantity is zero for any mode with non-zero energy:
$\overline{\phi_{k}(\beta)}=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}\phi_{k}(t,\beta)dt=0\,.$ (16)
This is dephasing: while the global pure state retains its coherence, any local measurement of this coherence will average to zero over time due to the destructive interference of the many oscillating modes with incommensurate frequencies. However, the entanglement entropy is a non-local quantity that is sensitive to these oscillating phase relationships. The ”double plateau” dynamics are the macroscopic manifestation of the interference between all these microscopic oscillators.
The GGE describes the stationary state after all such transient coherences have dephased. It is therefore constructed only from the expectation values of the truly conserved quantities. The GGE density matrix is thus
$\rho_{GGE}=\frac{1}{Z_{GGE}}\exp\left(-\sum_{k}\lambda_{k}\hat{n}_{k}\right)\,$, (17)
[…]
$\frac{1}{e^{\lambda_{k}}+1}=n_{k}=\frac{1}{e^{\beta E(k)}+1}\,.$ (18)
@@
This immediately yields the solution for the Lagrange multipliers: $\lambda_{k}=\beta E(k)$.
Substituting this result back into the definition of the GGE is the final step. The stationary state is:
$\rho_{\text{st}}=\frac{e^{-\beta H}}{Z}\,.$ (19)
Thus, the system thermalizes through a two-step process: first, the unitary evolution dephases the initial non-conserved quantum coherences (the pairing); Second, the remaining conserved populations guide the system to a GGE that, due to the special initial state, is identical to the Gibbs ensemble.
@@
Conclusion and outlooks — In this work, we study the problem of entanglement evolution after a quench from a TPQ state and find that the system eventually thermalize. We provide a complete theoretical explanation for this phenomenon. We prove that the final stationary state is the canonical Gibbs ensemble. This anomalous thermalization occurs because the special symmetries of the initial TPQ state force the expectation values of all the higher conserved charges that would lead to a GGE to be identically zero. The non-trivial dynamics are the macroscopic signature of the underlying microscopic mechanism. The initial state possesses non-zero anomalous pairing correlations, $\braket{c_{k}c_{-k}}\neq 0$, a form of quantum coherence forbidden in the final thermal state. The ”double plateau” is the real-time fingerprint of the coherent dephasing of these pairing terms, a process driven by the ballistic transport of stable quasiparticles. This work presents a rare exactly solvable model in which an effective thermal state emerges from the coherent evolution of a pure state, showing that the route to equilibrium in an integrable system can be remarkably rich and structured.
We find that the system thermalizes not because of chaos, but because the initial TPQ state is a highly non-generic pure state whose conserved quantities are already distributed in a perfectly thermal way. The non-trivial dynamics are the signature of the dephasing process by which the system sheds its initial quantum coherence to arrive at this final, incoherent thermal state.
Our results pave the way for exploring nonequilibrium dynamics from highly structured initial states and for extending the CFT quench framework to non-orientable manifolds. This perspective can be naturally generalized to richer scenarios, including inhomogeneous quenches such as the Möbius or sine-square deformation (SSD) [34, 35, 36, 37, 38, 39]. In addition, our analytical and numerical approaches provide a versatile tool for tracking more refined probes of quantum correlations, including mutual information and entanglement negativity.
@@ B.1 The imaginary-time evolution
$\frac{d\Gamma_{mn}(\tau)}{d\tau}=-\frac{i}{2}\text{Tr}\left(\rho(\tau)\{[\gamma_{m},\gamma_{n}],H\}\right)+2\Gamma_{mn}\text{Tr}(\rho(\tau)H)\,.$ (S41)
@@ D.2 The parent Hamiltonian of the crosscap states
This gives the pairing potential $\Delta_{k}=i\sin(kN)=i(-1)^{m}$.

### v2

@@ Abstract
We present an exact solution for the real-time dynamics following a quench from a thermal pure quantum (TPQ) state in an integrable system. Although equilibration is expected in this setting because the TPQ state already encodes the thermal expectation values of all conserved quantities, the approach to equilibrium shows highly nontrivial coherent dynamics. In the spin-1/2 XX chain, local observables become stationary at their thermal values after dephasing, while the entanglement entropy exhibits a characteristic double-plateau structure. We obtain this behavior exactly using three complementary approaches: two-dimensional (2D) conformal field theory (CFT) on the Klein bottle, an exact numerical evolution based on the matrix Riccati equation, and an asymptotically exact quasiparticle picture. These results demonstrate that the non-monotonic entanglement evolution is a macroscopic manifestation of the coherent dephasing of anomalous pairing correlations in the initial TPQ state.
@@
A central pillar of modern statistical mechanics is the eigenstate thermalization hypothesis (ETH), which provides a framework for understanding how isolated, chaotic quantum systems approach equilibrium [1, 2, 3, 4]. The ETH asserts that individual energy eigenstates of chaotic systems are already locally thermal and reflect canonical typicality, meaning that a generic pure state from a microcanonical energy shell is locally indistinguishable from a thermal ensemble [5]. Consequently, the real-time evolution from a state that is already thermal-like, such as a thermal pure quantum (TPQ) state [6], is expected to display little or no dynamical change in local observables or entanglement entropy. In contrast, integrable systems, endowed with an extensive number of conservation laws, do not follow ETH and retain detailed memory of their initial conditions, relaxing instead to nontrivial stationary states described by a generalized Gibbs ensemble (GGE) [7, 8].
@@
This raises a fundamental question: how does an integrable system evolve if it is prepared in a TPQ state? Does the dynamics retain the imprint of integrability, or reflect the thermal nature of the initial state? In this Letter, we provide an exact solution to this question and uncover a third dynamical paradigm, which is neither the trivial evolution of a chaotic system nor the standard non-thermal relaxation to a GGE. We consider a special TPQ state $|\Psi_{\beta}\rangle$ in the integrable spin-1/2 XX chain, with both the preparation and time evolution governed by the same Hamiltonian $H$.The TPQ state is constructed by imaginary-time evolving a nonlocal, maximally entangled crosscap state $|\mathcal{C}\rangle$ [9, 10, 11, 12, 13, 14, 15, 16, 17] as $|\Psi_{\beta}\rangle\propto e^{-\beta H/4}|\mathcal{C}\rangle$ [19, 20, 18].
The resulting dynamics reveal a distinctive, highly coherent route to equilibrium. Local observables reach to the thermal value after dephasing, yet the entanglement entropy displays a nonmonotonic evolution with a universal “double-plateau” structure. We establish this behavior from three complementary and exact perspectives. First, using 2D conformal field theory (CFT), we derive an analytical formula for the time-dependent entanglement entropy in terms of elliptic theta functions by evaluating vertex-operator correlators on the Klein bottle geometry. Second, we obtain an exact numerical benchmark by solving the governing matrix Riccati equation for the TPQ covariance matrix and evolving it in real time. Third, we show that the entire profile is quantitatively captured by an asymptotically exact quasiparticle picture [21, 22, 23, 24, 25], where the ballistic propagation of antipodally entangled quasiparticle pairs governs the coherent dephasing underlying the entropy double plateau. The perfect agreement among these approaches demonstrates that the phenomenon is exact and universal.
While it has been established that isolated integrable systems generally relax to non-thermal GGEs [26, 27], exceptions of true Gibbs-like equilibration are extremely rare and typically restricted to CFT or infinite-temperature cases [28, 29]. Our results provide an explicit lattice realization of a pure-state evolution that attains a canonical Gibbs ensemble, showing that equilibrium in integrable systems can emerge through a coherent, structured dephasing process rather than through chaos. This defines a new, exactly solvable pathway to quantum equilibration.
@@
Thermal Pure Quantum State — In 2D CFT, a crosscap state $|\mathcal{C}\rangle$ is a boundary state [30] embodying a non-orientable spacetime, defined by the constraint $(L_{n}-(-1)^{n}\bar{L}_{-n})|\mathcal{C}\rangle=0$ on the Virasoro generators [9]. While abstract, its physical essence is captured on a lattice of length $L$ as an entangled antipodal pair (EAP) state [18]. For a spin-1/2 chain, this is
@@
The Numerical Benchmark — To validate our analytical CFT predictions, we perform exact numerical simulations on the lattice realization of the theory, namely the spin-1/2 XX chain at half filling. The Hamiltonian for a system of $L=2N$ sites with periodic boundary conditions is
@@
After a Jordan-Wigner transformation, $\sigma_{j}^{+}=\left(\prod_{k=1}^{j-1}(-\sigma_{k}^{z})\right)c_{j}^{\dagger}$, $\sigma_{j}^{-}=\left(\prod_{k=1}^{j-1}(-\sigma_{k}^{z})\right)c_{j},\sigma_{j}^{z}=2c_{j}^{\dagger}c_{j}-1$. The model is described by a Hamiltonian of non-interacting fermions, $H=-2J\sum_{j=1}^{2N}(c_{j}^{\dagger}c_{j+1}+\text{h.c.})$, with an anti-periodic boundary condition. The single-particle energy dispersion is given by $E(k)=-4J\cos(k)$. The ground state is formed by filling all negative energy states, which corresponds to half-filling ($N_{f}=N$). The low-energy excitations occur near the two Fermi points, $k_{F}=\pm\pi/2$, where the dispersion becomes linear: $E(k_{F}\pm q)\approx\pm(4J)q$. This linear dispersion is the defining characteristic of the massless (1+1)D Dirac fermion, with an emergent ”speed of light” (Fermi velocity) of $v_{F}=4J$ (in units where the lattice spacing $a=1$).
@@
To validate our analytical CFT predictions, we perform exact numerical simulations on the spin-1/2 XX chain. The initial TPQ state $|\Psi_{\beta}\rangle$ is prepared by evolving the crosscap state $|\mathcal{C}\rangle$ in imaginary time, as described in Eq. (2). For the numerical implementation, we employ a fermionic representation of the system via the Jordan-Wigner transformation. It is important to note a subtle but important technical point regarding the lattice crosscap state itself. As was recently pointed out in Ref. [15], the Jordan-Wigner transformation of the strict spin crosscap state of Eq. (1) results in a superposition of two distinct fermionic Gaussian states. For simplicity and to maintain consistency with the Gaussian state formalism used in our analytical and quasiparticle pictures, our numerical simulation, following the approach in Ref. [16], targets the evolution from a Gaussian version of the crosscap state $|\mathcal{C}\rangle=\bigotimes_{j=1}^{L/2}\frac{1}{\sqrt{2}}(1+c_{j}^{\dagger}c_{j+L/2}^{\dagger})|0\rangle$. We expect that this simplification does not affect the qualitative features of the entanglement dynamics reported here, as the essential non-local entanglement structure is captured by either Gaussian component, though a quantitative investigation of the full superposition remains an interesting direction for future work.
@@
Figure 1: Comparison between the exact numerical results and the CFT prediction for the entanglement entropy $S_{A}(t)$. Solid lines: numerical data; circles: CFT prediction. Left: $L=500,l=80,\beta=20,v_{F}=1$. Right: $L=400,l=100,\beta=40,v_{F}=1$.
As shown in Fig. 1, the CFT prediction (symbols) is in excellent quantitative agreement with the exact lattice calculation (solid line), accurately capturing both the double-plateau structure and the characteristic time scales after the rescaling $\beta\to 2\pi v_{F}\beta/L,\sigma\to 2\pi l/L,t\to 2\pi v_{F}t/L$ in Eq. (4).
@@
Figure 2: Comparison between the exact numerical results and the quasiparticle prediction for the entanglement entropy $S_{A}(t)$. Solid lines: numerical data; dashed line: quasi-particle prediction. Left: $L=600,l=200,\beta=1,J=0.5$. Right: $L=1000,l=400,\beta=4,J=0.5$.
@@
Mechanism of Coherent Equilibration — We now elucidate the microscopic mechanism underlying the observed equilibration. The process is not driven by chaos but by the dephasing of the initial state’s quantum coherences, which can be described exactly in our model. After dephasing, the stationary state coincides with a GGE that, for this particular initial condition, reduces identically to the canonical Gibbs ensemble.
A defining feature of the TPQ state $|\Psi_{\beta}\rangle$ is its nonzero anomalous pairing amplitude,
$\phi_{k}(\beta)=\langle c_{k}c_{-k}\rangle=\frac{-i\sin(kN)}{2\cosh[E(k)\beta/2]}\,$, (15)
which signals quantum coherence between sectors of different fermion number. In the thermal Gibbs ensemble, such coherence vanishes identically. To understand how this coherence disappears, we analyze the BCS representation of the TPQ state in the thermodynamic limit,
$|\Psi_{\beta}(t)\rangle=\prod_{k>0}\!\left[u_{k}(\beta)+v_{k}(\beta)e^{-2iE(k)t}c_{k}^{\dagger}c_{-k}^{\dagger}\right]\!|0\rangle\,.$ (16)
The occupations $n_{k}=\langle c_{k}^{\dagger}c_{k}\rangle$ remain constant, while the pairing amplitude oscillates as $\phi_{k}(t,\beta)=\phi_{k}(0,\beta)e^{-i2E(k)t}$. Hence, observables sensitive to this coherence, such as $\langle c_{j}c_{j+1}\rangle$, display explicit temporal oscillations,
$\langle c_{j}c_{j+1}\rangle_{t}=\frac{1}{L}\sum_{k}e^{-ik(2j+1)}\phi_{k}(0)e^{-2iE(k)t}\,$, (17)
showing that the initial state is not stationary.
At long times, these oscillations dephase, and the time average satisfies $\overline{\phi_{k}(t)}=0$ for all $E(k)\neq 0$. Consequently, $\overline{\langle c_{j}c_{j+1}\rangle_{t}}=0$, exactly matching the thermal expectation value $\mathrm{Tr}(\rho_{\mathrm{th}}c_{j}c_{j+1})=0$. In contrast, number-conserving observables such as $\langle c_{j}^{\dagger}c_{j+1}\rangle$ remain stationary and coincide with their thermal values at all times. Thus, all local observables approach the canonical values once the coherence encoded in $\phi_{k}$ has dephased. The entanglement entropy provides the macroscopic witness of this microscopic process. The nonmonotonic “double-plateau” profile directly tracks the coherent loss and reorganization of pairing correlations.
After dephasing, the remaining stationary state is fully characterized by the conserved occupations $n_{k}$. The corresponding GGE density matrix reads
$\rho_{\mathrm{GGE}}=\frac{1}{Z_{\mathrm{GGE}}}\exp\!\left[-\sum_{k}\lambda_{k}\hat{n}_{k}\right]\,$, (18)
@@
with Lagrange multipliers fixed by $\mathrm{Tr}(\rho_{\mathrm{GGE}}\hat{n}_{k})=n_{k}=(e^{\beta E(k)}+1)^{-1}$, which yields $\lambda_{k}=\beta E(k)$. Substituting this into $\rho_{\mathrm{GGE}}$ gives
$\rho_{\mathrm{st}}=\frac{e^{-\beta H}}{Z}.$ (19)
Therefore, equilibration proceeds in two coherent stages: (i) the unitary dynamics dephase the nonconserved pairing coherences, and (ii) the remaining conserved populations define a GGE that, for this special state, coincides exactly with the Gibbs ensemble. This dephasing-based mechanism provides an explicit, solvable realization of thermalization without chaos.
@@
Conclusion and outlooks — In this work, we study the entanglement evolution following a quench from a TPQ state and uncover a distinctive, coherent pathway to equilibrium. We provide a complete theoretical account of this process and show that the stationary state coincides with the canonical Gibbs ensemble. This apparent thermal behavior arises because the special symmetries of the initial TPQ state enforce all higher conserved charges that would otherwise lead to a GGE to vanish identically. The resulting nontrivial dynamics are the macroscopic signature of an underlying microscopic mechanism. The initial state possesses nonzero anomalous pairing correlations, $\langle c_{k}c_{-k}\rangle\neq 0$, representing quantum coherence forbidden in the final thermal state. The observed “double plateau” is the real-time fingerprint of the coherent dephasing of these pairing correlations, a process driven by the ballistic propagation of stable quasiparticles. This highlights a central finding of our work that a structured, nonmonotonic entanglement evolution can emerge as the macroscopic manifestation of microscopic quantum coherence during equilibration. Conceptually, this connects our results to other observations of intermediate-time coherent dynamics in the XX chain, such as the emergence and disappearance of short-range coherence in quenches leading to an infinite-temperature state [40]. While the initial states and final ensembles differ, both reveal that the route to equilibrium in integrable systems can be remarkably rich and structured.
Our results pave the way for exploring nonequilibrium dynamics from highly structured initial states and for extending the CFT quench framework to non-orientable manifolds. The CFT analysis and exact Gaussian numerics apply directly to mutual information and entanglement negativity, while the quasiparticle picture can be adapted with minor modifications. Exploring interacting integrable models in this context would be both interesting and important for future work. This perspective can be generalized to richer scenarios, including inhomogeneous quenches such as the Möbius or sine-square deformation (SSD) [41, 42, 43, 44, 45, 46]. The experimental exploration of crosscap-type initial states and the associated entanglement dynamics would be a fascinating direction for future studies, potentially realizable in quantum-simulation platforms such as ultracold atoms or superconducting qubits.
@@ B.1 The imaginary-time evolution
$\frac{d\Gamma_{mn}(\tau)}{d\tau}=-\frac{i}{2}\text{Tr}\left(\rho(\tau)\{[\gamma_{m},\gamma_{n}],H\}\right)+2\Gamma_{mn}(\tau)\text{Tr}(\rho(\tau)H)\,.$ (S41)
@@ D.2 The parent Hamiltonian of the crosscap states
This gives the pairing potential $\Delta_{k}^{0}=i\sin(kN)=i(-1)^{m}$.